OpenStudy (anonymous):

-5x-2y=-20 2x+8y=8 is the answer (-3,4), (-3,-4), (4,0) or (4,3)?

OpenStudy (anonymous):

We want to determine -5x-2y=-20 and 2x+8y=8 Simplify the equation for line 1 to get it into our y = mx + b format: --5x - 2y = -20 -2y = -5x - 20 Divide each side of the equation by -2 to isolate y: -2y -2 = -5x - 20 -2 Simplifying and evaluating, we have: y = -2.5x + 10 Therefore, the slope of line equation 1 = -2.5 Simplify the equation for line 2 to get it into our y = mx + b format: 2x + 8y = 8 8y = -2x + 8 Divide each side of the equation by 8 to isolate y: 8y 8 = -2x + 8 8 Simplifying and evaluating, we have: y = -0.25x + 1 Therefore, the slope of line equation 2 = -0.25 Since y = -2.5x + 10 and y = -0.25x + 1, set each line equation equal to each other and solve for x: -2.5x + 10 = -0.25x + 1 -2.5x - -0.25x = 1 - 10 - 2.25x = -9 x = -9/ - 2.25 x = 4 Now that we have x, plug it into equation 1 to find y y = -2.5 * (4) + 10 y = -10 + 10 y = 0 Therefore, our intersection point = (4, 0) Calculate the product of the 2 slopes: Slope 1 * Slope 2 = -2.5 * -0.25 = 0.625 Since the product of the 2 slopes <> - 1, the lines are not perpendicular The 2 lines intersect at (4, 0)